OpenStudy (anonymous):

(x-5) (4x-38) find the length

OpenStudy (freckles):

I kinda don't understand.

OpenStudy (freckles):

Can you give more info?

OpenStudy (anonymous):



OpenStudy (freckles):

is D the midpoint?

OpenStudy (freckles):

like does it say D is the midpoint?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

that would mean ED=DG

OpenStudy (freckles):

so that would mean you need to solve x-5=4x-38 first

OpenStudy (anonymous):

implifying X + -5 = 4x + -38 Reorder the terms: -5 + X = 4x + -38 Reorder the terms: -5 + X = -38 + 4x Solving -5 + X = -38 + 4x Solving for variable 'X'. Move all terms containing X to the left, all other terms to the right. Add '5' to each side of the equation. -5 + 5 + X = -38 + 5 + 4x Combine like terms: -5 + 5 = 0 0 + X = -38 + 5 + 4x X = -38 + 5 + 4x Combine like terms: -38 + 5 = -33 X = -33 + 4x Simplifying X = -33 + 4x

OpenStudy (freckles):

but we have to one step further after finding x

OpenStudy (anonymous):

huh?

OpenStudy (freckles):

because it didn't ask for x it asked for the length of EG

OpenStudy (freckles):

but finding x will help us to do that

OpenStudy (anonymous):

how u set it up

OpenStudy (freckles):

well I was saying ED=DG since D is the midpoint

OpenStudy (freckles):

so that means we have x-5=4x-38

OpenStudy (freckles):

we solve this for x

OpenStudy (freckles):

then since EG=2(x-5) or 2(4x-38) or even (x-5)+(4x-38) you can plug in one of these 3 expressions to find the length of EG

OpenStudy (freckles):

all of them will give you the same number

OpenStudy (freckles):

so you only need one of those expressions for EG

OpenStudy (freckles):

let's do the first thing solve x-5=4x-38 first

OpenStudy (anonymous):

how u solve if x don't have a number

OpenStudy (freckles):

x is a number such that x-5=4x-38

OpenStudy (freckles):

get all your like terms on separate sides

OpenStudy (freckles):

get your x's on one side and you constant values on the other side

OpenStudy (anonymous):

can you show me

OpenStudy (freckles):

x-5=4x-38 x and 4x on separate sides they need to be on the same side so you to subtract one of them on both sides try

OpenStudy (anonymous):

3x=-5-38?